OpenStudy (anguyennn):

how do you simplify -1 1/3-5/6+7/4-2

OpenStudy (anguyennn):

\[-1\frac{ 1 }{ 3 }-\frac{ 5 }{ 6 }+\frac{ 7 }{ 4 }-2\]

OpenStudy (anonymous):

Try simplifying the fractions first.

OpenStudy (anonymous):

This is both mixed numbers and unequal fractions, so decide which you're better with and convert to that.

OpenStudy (anguyennn):

how?

OpenStudy (anguyennn):

I moved the -2 over with -1

OpenStudy (anguyennn):

and then I multiplied the denominators together so that they can all equal to 12

OpenStudy (anonymous):